If tan θ = - \sqrt(3)/3 and cos θ > 0; Find csc θ.


1
Expert's answer
2021-11-25T23:38:59-0500

let "\theta" = k

Now, tan k = - ("\sqrt3" / 3) = - (1/"\sqrt3" )

cos k > 0


Now tan k is negative when k lies in second or fourth quadrant. And cos k is positive in first and fourth quadrant.

So we can infer that k lies in the fourth quadrant.


Now tan k = - (1/"\sqrt3" ) => k = - 30"\degree" => k = (360 - 30)"\degree" = 330"\degree"

So k = 330"\degree"


Now cosec k = cosec 330"\degree" = cosec (360 - 30)"\degree" = - cosec 30"\degree" = -2


therefore the answer is -2.